Refer to the polynomial f(x)=3x(x−3)2(x+2)3(x2−3)
This function has five zeroes.
1. Determine the degree of this polynomial.
2. State the five zeroes.
3. State the multiplicity of each zero. Make it clear which multiplicity corresponds to which zero.

Answer by stanbon(75887)

Refer to the polynomial f(x)=3x(x−3)^2(x+2)^3(x^2−3)
This function has five zeroes.
1. Determine the degree of this polynomial.
Ans: 8
============================
2. State the five zeroes.
Ans: x = 0,3,-2,sqrt(3),-sqrt(3)
------------------------------------------
3. State the multiplicity of each zero. Make it clear which multiplicity corresponds to which zero.
x = 0;; multi = 1
X = 3;; multi = 2
x = -2;; multi = 3
x = sqrt(3);; multi = 1
x = -sqrt(3);; multi = 1
===============================
Cheers,
Stan H.
-------------------------------